OpenStudy (bbb911):

This is Geometry: Write an equation of a circle with center (6,-9) and diameter= 2 sqrt of 7

OpenStudy (anonymous):

So this is just like the last one you did except you have the diameter instead of the radius. Can you find the radius using the diameter?

OpenStudy (bbb911):

no im stupped :(

OpenStudy (anonymous):

The diameter is two times the radius.

OpenStudy (anonymous):

You aren't stupid. You just have to try.

OpenStudy (bbb911):

i am in math i always fail :( in english im really good at it

OpenStudy (anonymous):

If you fail it is because you aren't trying. So if the diameter is 2 times the radius, how can you find the radius?

OpenStudy (bbb911):

even if i try :( did that once ended up gettting a D :(

OpenStudy (anonymous):

You aren't dumb.

OpenStudy (bbb911):

:(

OpenStudy (anonymous):

So if the diameter is 2 times the radius, and the diameter is 2 times sqrt(7) what is the radius?

OpenStudy (bbb911):

4?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

\[D = 2(r)\]\[D = 2(\sqrt{7})\] What is r?

OpenStudy (bbb911):

6

OpenStudy (anonymous):

Look at what I wrote.. What is different about those two lines?

OpenStudy (anonymous):

The only difference is that one had an r and one had a \(\sqrt{7}\)

OpenStudy (anonymous):

Doesn't it seem likely that r is \(\sqrt{7}\)

OpenStudy (bbb911):

ohhhhhh ok so the radius is 7

OpenStudy (anonymous):

No, the radius is the square root of 7.

OpenStudy (bbb911):

say its sq 8 it would the r would be sqrt of 8 to?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

D = 2(r) If someone tells you that D is 2\(\sqrt{8}\) Then you know that r is\(\sqrt{8}\)

OpenStudy (bbb911):

oh i get it now

OpenStudy (bbb911):

oh ok sounds good

OpenStudy (anonymous):

If I tell you D is 10, then 2r = 10 so what is r?

OpenStudy (bbb911):

10

OpenStudy (anonymous):

no

OpenStudy (anonymous):

D = 10

OpenStudy (anonymous):

2r = 10

OpenStudy (anonymous):

r = ?

OpenStudy (bbb911):

2

OpenStudy (anonymous):

If 10 is 2 times r, then isn't r half of 10?

OpenStudy (bbb911):

half of 10 ?

OpenStudy (anonymous):

Ok, lets look at it this way. If I tell you I'm thinking of a number. And I won't tell you what that number is.. We'll call it r.

OpenStudy (bbb911):

oh ok

OpenStudy (anonymous):

Now I tell you that if you take my number 'r' and multiply it by 2 you will get 12. Can you tell me what the number I'm thinking of is?

OpenStudy (bbb911):

10

OpenStudy (anonymous):

2 times 10 is 12?!?!

OpenStudy (bbb911):

ooo wait i messed up

OpenStudy (bbb911):

its 6

OpenStudy (anonymous):

Right! Are you a psychic!?!? How did you know?

OpenStudy (bbb911):

i just times 6 and 2 hehe

OpenStudy (anonymous):

That's a hard way. You have to guess a lot. Couldn't you just divide 12 by 2?

OpenStudy (bbb911):

yea so true

OpenStudy (anonymous):

Ok. So lets play the guessing game again. I'm thinking of a radius. I wont tell you ! It's a secret! Shh.. But I will tell you that the diameter is 64. And that diameter is 2 times my radius. What is my radius?

OpenStudy (bbb911):

32

OpenStudy (anonymous):

OMG you are a psychic! I knew it!

OpenStudy (bbb911):

yaya:D

OpenStudy (anonymous):

Just to prove your mental powers we'll do one final one. I'm thinking of a radius. The diameter (2 times the radius) is \(4\sqrt{3}\). What is the radius?

OpenStudy (bbb911):

okay thats a bit harderder

OpenStudy (anonymous):

But not too hard.

OpenStudy (anonymous):

What does \(4\sqrt{3}\) mean?

OpenStudy (bbb911):

the r is 3

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

It's a bit tricky but I think you can do it.

OpenStudy (anonymous):

What does \(4\sqrt{3}\) mean?

OpenStudy (bbb911):

4 sqrt of 3

OpenStudy (anonymous):

Right. But that _means_ you have four \(\sqrt{3}\) added together. It also _means_ \(4 \times \sqrt{3}\)

OpenStudy (anonymous):

So can you figure out what: \[4\times \sqrt{3}\div 2 =\ ?\]

OpenStudy (anonymous):

\[4\times \sqrt{3}\over 2\]

OpenStudy (bbb911):

that would be 6 then

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

\[{4(\sqrt{3}) \over 2} = {4 \over 2}(\sqrt{3})\]

OpenStudy (anonymous):

What is 4/2?

OpenStudy (bbb911):

oh 2

OpenStudy (anonymous):

So what is \[{4 \over 2}(\sqrt{3})\]

OpenStudy (anonymous):

Taking too long.. You must have gotten confused. If 4/2 IS 2 Then (4/2) times sqrt(3) IS 2 times sqrt(3)

OpenStudy (bbb911):

umm a lil

OpenStudy (anonymous):

Read what I wrote. Out loud if needed.

OpenStudy (anonymous):

Sometimes it helps to hear things than to see them.

OpenStudy (bbb911):

oh both 4 and 2 time sq 3

OpenStudy (anonymous):

No.

OpenStudy (bbb911):

then im confused

OpenStudy (anonymous):

4 divided by 2 IS 2. 6 divided by 2 IS 3. (4 divided by 2) times (sqrt 5) IS 2 times (sqrt 5) (6 divided by 2) times (sqrt 27) IS 3 times (sqrt 27)

OpenStudy (anonymous):

If 4 divided by 2 is exactly the same thing as 2 then any time you see 4 divided by 2 you can just write a 2 there instead.

OpenStudy (bbb911):

oh i think answer should be 28=(x+6)sq + (y-9)sq am i correct?

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

Close

OpenStudy (bbb911):

does the 6 have to be negative and the nine positive?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Also your radius is wrong.

OpenStudy (bbb911):

ok is the answer 7=(x-6) sq + (y+9)sq

OpenStudy (anonymous):

Yes

OpenStudy (bbb911):

yaya i did it :3

OpenStudy (anonymous):

=)

OpenStudy (bbb911):

thanks so much for your help

OpenStudy (anonymous):

Of course. I suggest that if you get some free time you practice some simple equations. I think if you invest a little time in the areas you need work you will find solving these other problems much much easier.

OpenStudy (bbb911):

yea thats true thanks for the advice its very helpful